OpenStudy (anonymous):

What integer needs to be added to the expression to create a perfect square? x2 + 14x

OpenStudy (luigi0210):

take the second part of 14x and divide it by two

OpenStudy (anonymous):

that equals 7

OpenStudy (luigi0210):

Yup! so use that in (x+a)(x+a)

OpenStudy (luigi0210):

So it's (x+7)(x+7),when you multiply it put it gives you a x^2 and a 14x,and the 7*7=49

OpenStudy (anonymous):

so x^2 + 14x + 49

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

A. -3, 5 B. 3, -5 C. 1, -15 D. 3, 5 these are the answers I don't want the answer I want to figure out how to get it

OpenStudy (anonymous):

sorry wrong question these are the answers A. 7 B. 14 C. 49 D. 196

OpenStudy (luigi0210):

We already solved, the integer is 49

OpenStudy (luigi0210):

because of the 14/2=7 and 7*7=49

OpenStudy (anonymous):

oh ok thank you

OpenStudy (luigi0210):

It's asking for the last number in the equation and that would be 49

OpenStudy (luigi0210):

And you are welcome

OpenStudy (anonymous):

thanks